# Non-Gaussianities in primordial black hole formation and induced gravitational waves

### v1

@@ Abstract
The most promising mechanism of generating primordial black holes (PBHs) is by the enhancement of power spectrum of the primordial curvature perturbation, which is usually accompanied by the the enhancement of non-Gaussianity that crucially changes the abundance of PBHs. In this review I will discuss how non-Gaussianity is generated in single field inflation as well as in the curvaton scenario, and then discuss how to calculate PBH mass function with such non-Gaussianities. I also show non-Gaussianity only has mild effects on the induced gravitational waves (GWs), which gives robust predictions in the mHz and nHz GW experiments.
@@ I Overview
A related phenomenon is the stochastic gravitational wave background induced by the enhanced curvature perturbation. In General Relativity, the curvature perturbation $\mathcal{R}$ and tensor perturbation $h_{ij}$ are only coupled at nonlinear order, and the most important coupling is the tensor-scalar-scalar type $\sim h_{ij}\partial_{i}\mathcal{R}\partial_{j}\mathcal{R}$ Matarrese et al. (1993, 1994, 1998); Noh and Hwang (2004); Carbone and Matarrese (2005); Nakamura (2007); Ananda et al. (2007); Osano et al. (2007); Baumann et al. (2007). Therefore, the curvature perturbation can source the GWs after it reenters the Hubble horizon, which is called scalar induced secondary GWs. The peak frequency, as well as the PBH mass generated by such a perturbation, are determined by the horizon scale at the reentry moment, which are connected by Saito and Yokoyama (2009)
@@
At the peak scale, the spectral energy density of the induced GWs we observe today (normalized by the critical density), is roughly $\Omega_{\text{GW,0}}\equiv\rho_{\mathrm{cr}}^{-1}\left(\mathrm{d}\rho_{\mathrm{GW}}/\mathrm{d}\ln k\right)_{t_{0}}\sim 10^{-6}\mathcal{P}_{\mathcal{R}}^{2}$, where $\mathcal{P}_{\mathcal{R}}$ is the primordial power spectrum of $\mathcal{R}$. The nearly scale-invariant power spectrum extrapolated from CMB scales predicts a scale-invariant induced GW with an amplitude of $\Omega_{\text{GW,0}}\sim 10^{-24}$, far less than observational bound Baumann et al. (2007). However, for the enhanced curvature perturbation with abundant PBH formation ($\mathcal{P_{R}}\sim\mathcal{O}(10^{-2})$), the induced GW is also enhanced to $\Omega_{\text{GW,0}}\sim 10^{-10}$, which reaches the detectable range of many experiments like the next-generation ground- and space-borne interferometers and pulsar timing arrays. As both $\Omega_{\text{GW,0}}$ and PBH abundance $f_{\mathrm{PBH}}$ depend on $\mathcal{P_{R}}$, they can be cross-checked to probe the primordial power spectrum on small scales Saito and Yokoyama (2009, 2010); Bugaev and Klimai (2010); Assadullahi and Wands (2010); Bugaev and Klimai (2011). For more discussion on the induced GWs, see Domènech (2024).
However, to enhance the power spectrum of the curvature perturbation, usually the slow-roll attractor is deviated, while the shape and amplitude of the enhanced power spectrum, as well as the statistics are also changed, depending on the model. Even a small non-Gaussianity can change the relation between the PBH abundance and induced GW spectrum significantly, as the induced GW depends mainly on the variance of the PDF of the curvature perturbation, while the PBH abundance depends on the high-$\sigma$ tail of the PDF. A small change in the PDF, i.e. deviation of the Gaussianity, can give rise to a huge enhancement/suppression in the PBH abundance Bullock and Primack (1997); Ivanov (1998); Yokoyama (1998); Pina Avelino (2005); Seery and Hidalgo (2006); Hidalgo (2007); Byrnes et al. (2012); Young and Byrnes (2013); Young et al. (2016); Atal and Germani (2019); Yoo et al. (2019); Kehagias et al. (2019); Mahbub (2020); Riccardi et al. (2021); Taoso and Urbano (2021); Biagetti et al. (2021); Kitajima et al. (2021); Young (2022); Escrivà et al. (2022a); Matsubara and Sasaki (2022); Gow et al. (2022), but only change the amplitude of the induced GW spectrum mildly Nakama et al. (2017); Garcia-Bellido et al. (2017); Cai et al. (2019a); Unal (2019); Unal et al. (2020); Adshead et al. (2021); Garcia-Saenz et al. (2022); Abe et al. (2022).
@@
which was used to describe the curvaure perturbation on small scales for PBH formation Byrnes et al. (2012); Young and Byrnes (2013); Tada and Yokoyama (2015); Young and Byrnes (2015); Young et al. (2016); Franciolini et al. (2018); Ando et al. (2018a); Atal and Germani (2019); Passaglia et al. (2019) as well as the induced GWs Nakama et al. (2017); Garcia-Bellido et al. (2017); Cai et al. (2019a); Unal (2019); Unal et al. (2020). However, in many models, the non-linear parameter $f_{\mathrm{NL}}$ reaches $\mathcal{O}(1)$ Namjoo et al. (2013); Martin et al. (2013); Chen et al. (2013); Motohashi et al. (2015); Davies et al. (2022); Namjoo (2023); Namjoo and Nikbakht (2024), so the higher orders in (2) might be important, which are highly model-dependent. These higher order terms must be taken into account in the PBH formation. Fortunately, by $\delta N$ formalism, a fully nonlinear curvature perturbation can be derived for single field inflation and for the curvaton scenario, which indeed alters the PDF of $\mathcal{R}$ completely in large $\mathcal{R}$ regime. Various methods are developed to calculate the PBH abundance for such a PDF, including the Press-Schechter-type formalism Biagetti et al. (2021); Gow et al. (2022); Ferrante et al. (2022) and the peak theory Green et al. (2004); Yoo et al. (2018); Germani and Musco (2019); Atal et al. (2019, 2020a); Young and Musso (2020); Yoo et al. (2021); Taoso and Urbano (2021); Riccardi et al. (2021); Kitajima et al. (2021); Young (2022). The Press-Schechter-type formalism for Gaussian curvature perturbation is straightforward to be further extended to non-Gaussian case.
@@
Considering non-Gaussianity on the PBH formation and induced GW generation is important in many theoretical issues and observational problems in cosmology.
@@
  • PBHs might be the supermassive or stupendously large BHs which seed the galaxy or structure formation Bean and Magueijo (2002); Kawasaki et al. (2012); Carr and Silk (2018); Carr et al. (2021). To satisfy the stringent constraints from CMB $\mu$- and $y$-distortion, large non-Gaussianities must be included Nakama et al. (2016, 2018, 2019); Atal et al. (2020b); Carr and Silk (2018); Liu and Bromm (2022); Biagetti et al. (2023); Gouttenoire et al. (2023); Hooper et al. (2023).
@@
  • According to the current observational constraints, asteroid-mass PBH can be all the dark matter, of which the concomitant induced GWs are detectable by space-borne interferometers LISA Barausse et al. (2020); Arun et al. (2022); Auclair et al. (2023); Bagui et al. (2023), Taiji Ren et al. (2023), and TianQin Liang et al. (2021). As positive non-Gaussianity can enhance the PBH abundance greatly, the required power spectrum of the curvature perturbation for $\Omega_{\mathrm{PBH}}=\Omega_{\mathrm{DM}}$ is reduced with positive non-Gaussianity, and the induced GW spectrum is also reduced, yet it is still above the LISA/Taiji/TianQin sensitivity curve Cai et al. (2019a); Bartolo et al. (2019); Unal (2019). The large non-Gaussian limit with $f_{\mathrm{NL}}\gg 1$ (i.e. the $\chi^{2}$-distribution) can be realized in the curvaton model when curvaton is negligible at its decay Bartolo et al. (2004); Enqvist and Nurmi (2005); Sasaki et al. (2006); Pi and Sasaki (2023a); Ferrante et al. (2022).
@@
  • In 2023, the Pulsar Timing Array collaborations NANOGrav Agazie et al. (2023a, b), EPTA combined with InPTA Antoniadis et al. (2023a, b, c), PPTA Zic et al. (2023); Reardon et al. (2023a, b), CPTA Xu et al. (2023), as well as IPTA Agazie et al. (2023c), announced the strong evidence of nHz stochastic GW background, while the induced GW interpretation is favored Afzal et al. (2023). However, as the amplitude of the nHz GW spectrum is relatively high ($\sim 10^{-9}$), sub-solar-mass PBHs might be overproduced, i.e. $\Omega_{\mathrm{PBH}}>\Omega_{\mathrm{DM}}$ This tension can be solved by introducing a negative non-Gaussianity with $f_{\mathrm{NL}}\lesssim-1$ to suppress the PBH formation Franciolini et al. (2023a); Liu et al. (2023). Such a large and negative $f_{\mathrm{NL}}$ should be seen as the leading order of a fully nonlinear curvature perturbation, which can be realized in the curvaton scenario when the curvaton dominates at its decay moment Pi and Sasaki (2023a); Ferrante et al. (2022); Chen et al. (2023).
@@
  • It was shown that PBHs obey a Poisson distribution when the density perturbation is Gaussian Ali-Haïmoud (2018); Desjacques and Riotto (2018), while non-Gaussianity can make PBHs to cluster initially on small scales Tada and Yokoyama (2015); Young and Byrnes (2015); Franciolini et al. (2018); Young and Byrnes (2020); Suyama and Yokoyama (2019); De Luca et al. (2021), which enhances the merger rate of PBHs significantly Clesse and García-Bellido (2017); Raidal et al. (2017). Such clustered PBHs can escape from the microlensing detections, which assume Poisson distribution García-Bellido and Clesse (2018).
@@
  • When the power spectrum of the curvature perturbation is enhanced on small scales, the loop corrections on CMB scales might be large, which may jeopardize the mechanism of enhancing the power spectrum on small scales to generate PBHs Kristiano and Yokoyama (2022, 2023). It was soon realized that the loop correction can be suppressed when the transition to slow-roll inflation is smooth Riotto (2023); Firouzjahi (2023); Firouzjahi and Riotto (2024); Firouzjahi (2024a), as it is proportional to $f_{\mathrm{NL}}$. For recent discussions, see Motohashi and Tada (2023); Fumagalli (2023); Cheng et al. (2023); Franciolini et al. (2023b); Fumagalli et al. (2023); Maity et al. (2023); Iacconi et al. (2023); Firouzjahi (2024b).
@@
For more discussion on the observational implications, see Part V and the references therein, for instance the review papers Carr et al. (2020); Carr and Kuhnel (2020); Carr et al. (2024).
@@
This paper is organized as follows. In Section II we review calculation of PBH mass function, emphasizing on when and how the non-Gaussianity comes into play. In Section III, we study some typical inflation models which can generate large non-Gaussianity: the ultra-slow-roll inflation, constant-roll inflation with a bumpy potential, inflation with piecewise quadratic potential, and the curvaton scenario. After getting the non-Gaussian curvature perturbations in these models, we calculate the PBH mass function and abundance by the Press-Schechter-type formalism in Section IV, as well as the scalar-induced GWs in Section V. We conclude by discussing some interesting topics to be studied in the future in Section VI.
@@ II.1 PBH abundance in the Gaussian case
In this section we will briefly review how to calculate the PBH mass function and abundance by the Press-Schechter formalism, emphasizing on where and how the non-linearities and primordial non-Gaussianities enter. Our starting point is the power spectrum of the primordial curvature perturbation on comoving slices, $\mathcal{P_{R}}(k)$, generated at the end of inflation. The simplest Press-Schechter formalism uses the density contrast on comoving slices to calculate, which is related to $\mathcal{R}$ by the Poisson equation
$\delta\equiv\left(\frac{\delta\rho}{\rho_{0}}\right)_{c}=-\frac{2(1+w)}{5+3w}\frac{1}{H^{2}a^{2}}\nabla^{2}\mathcal{R}.$ (3)
@@
We will not discuss the issues of window function in this paper. The key point of the above calculation is that, if the curvature perturbation $\mathcal{R}$ is Gaussian, the density contrast is also Gaussian, guaranteed by the linear relation (3). Therefore the integration (6) is simply a complementary error function, which gives a Gaussian suppression $\sim\exp(-\delta^{2}/2\delta_{\mathrm{th}}^{2})$ at the large tail.

### v2

@@ Abstract
The most promising mechanism of generating PBHs is by the enhancement of power spectrum of the primordial curvature perturbation, which is usually accompanied by the the enhancement of non-Gaussianity that crucially changes the abundance of PBHs. In this review I will discuss how non-Gaussianity is generated in single field inflation as well as in the curvaton scenario, and then discuss how to calculate PBH mass function and induced gravitational waves (GWs) with such non-Gaussianities. When the PBH abundance is fixed, non-Gaussianity only has mild impact on the spectral shape of the induced GWs, which gives relatively robust predictions in the mHz and nHz GW experiments.
@@ I Overview
A related phenomenon is the stochastic gravitational wave background induced by the enhanced curvature perturbation. In General Relativity, the curvature perturbation $\mathcal{R}$ and tensor perturbation $h_{ij}$ are only coupled at nonlinear order, and the tensor-scalar-scalar type $\sim h_{ij}\partial_{i}\mathcal{R}\partial_{j}\mathcal{R}$ can induce second-order GWs Matarrese et al. (1993, 1994, 1998); Noh and Hwang (2004); Carbone and Matarrese (2005); Nakamura (2007); Ananda et al. (2007); Osano et al. (2007); Baumann et al. (2007)  [^2] For GWs induced by other couplings like scalar-tensor-tensor or tensor-tensor-tensor, see for instance Refs. Chang et al. (2023); Yu and Wang (2023); Bari et al. (2024); Picard and Malik (2023).. Therefore, the curvature perturbation can source the GWs after it reenters the Hubble horizon, which is called scalar induced secondary GWs. The peak frequency, as well as the PBH mass generated by such a perturbation, are determined by the horizon scale at the reentry moment, which are connected by Saito and Yokoyama (2009)
@@
At the peak scale, the spectral energy density of the induced GWs we observe today (normalized by the critical density), is roughly $\Omega_{\text{GW,0}}\equiv\rho_{\mathrm{cr}}^{-1}\left(\mathrm{d}\rho_{\mathrm{GW}}/\mathrm{d}\ln k\right)_{t_{0}}\sim 10^{-6}\mathcal{P}_{\mathcal{R}}^{2}$, where $\mathcal{P}_{\mathcal{R}}$ is the primordial power spectrum of $\mathcal{R}$. The nearly scale-invariant power spectrum extrapolated from CMB scales predicts a scale-invariant induced GW with an amplitude of $\Omega_{\text{GW,0}}\sim 10^{-24}$, far less than observational bound Baumann et al. (2007). However, for the enhanced curvature perturbation with abundant PBH formation ($\mathcal{P_{R}}\sim\mathcal{O}(10^{-2})$), the induced GW is also enhanced to $\Omega_{\text{GW,0}}\sim 10^{-10}$, which reaches the detectable range of many experiments like the next-generation ground- and space-borne interferometers and pulsar timing arrays. As both $\Omega_{\text{GW,0}}$ and PBH abundance $f_{\mathrm{PBH}}$ depend on $\mathcal{P_{R}}$, they can be cross-checked to probe the primordial power spectrum on small scales Saito and Yokoyama (2009, 2010); Bugaev and Klimai (2010); Assadullahi and Wands (2010); Bugaev and Klimai (2011). For more discussion on the induced GWs, see Chapter 18 Domènech (2024).
To enhance the power spectrum of the curvature perturbation, usually a deviation from the slow-roll attractor is required. Depending on the model, this not only changes the shape and amplitude of the enhanced power spectrum, but also the statistics of the curvature perturbation. Even a small non-Gaussianity can change the relation between the PBH abundance and induced GW spectrum significantly, as the induced GW depends mainly on the variance of the PDF of the curvature perturbation, while the PBH abundance depends on the high-$\sigma$ tail of the PDF. A small change in the PDF, i.e. deviation from the Gaussianity, can give rise to a huge enhancement/suppression in the PBH abundance Bullock and Primack (1997); Ivanov (1998); Yokoyama (1998); Pina Avelino (2005); Seery and Hidalgo (2006); Hidalgo (2007); Byrnes et al. (2012); Young and Byrnes (2013); Young et al. (2016); Atal and Germani (2019); Yoo et al. (2019); Kehagias et al. (2019); Mahbub (2020); Riccardi et al. (2021); Taoso and Urbano (2021); Biagetti et al. (2021); Kitajima et al. (2021); Young (2022); Escrivà et al. (2022a); Matsubara and Sasaki (2022); Gow et al. (2022), but only changes the amplitude of the induced GW spectrum mildly Nakama et al. (2017); Garcia-Bellido et al. (2017); Cai et al. (2019a); Unal (2019); Unal et al. (2020); Adshead et al. (2021); Garcia-Saenz et al. (2022); Abe et al. (2022).
@@
which is also used to describe the curvature perturbation on small scales for PBH formation Byrnes et al. (2012); Young and Byrnes (2013); Tada and Yokoyama (2015); Young and Byrnes (2015); Young et al. (2016); Franciolini et al. (2018); Ando et al. (2018a); Atal and Germani (2019); Passaglia et al. (2019); Özsoy and Tasinato (2023) as well as the induced GWs Nakama et al. (2017); Garcia-Bellido et al. (2017); Cai et al. (2019a); Unal (2019); Unal et al. (2020). However, in many models, the non-linear parameter $f_{\mathrm{NL}}$ reaches $\mathcal{O}(1)$ Namjoo et al. (2013); Martin et al. (2013); Chen et al. (2013); Motohashi et al. (2015); Davies et al. (2022); Namjoo (2023); Namjoo and Nikbakht (2024), so the higher orders in (2) might be important, which are highly model-dependent. These higher order terms must be taken into account in the PBH formation. Fortunately, by $\delta N$ formalism, a fully nonlinear curvature perturbation can be derived for single field inflation and for the curvaton scenario, of which the PDF of $\mathcal{R}$ can be calculated analytically. Various methods are developed to calculate the PBH abundance for such a PDF, including the Press-Schechter-type formalism Biagetti et al. (2021); Gow et al. (2022); Ferrante et al. (2022) and the theory of peaks Green et al. (2004); Yoo et al. (2018); Germani and Musco (2019); Atal et al. (2019, 2020a); Young and Musso (2020); Yoo et al. (2021); Taoso and Urbano (2021); Riccardi et al. (2021); Kitajima et al. (2021); Young (2022). The Press-Schechter-type formalism for Gaussian curvature perturbation is thoroughly reviewed in Chapter 7 Young (2024), which is straightforward to be further extended to non-Gaussian case.
@@
Considering non-Gaussianity in the PBH formation and induced GW generation is important in many theoretical issues and observational problems in cosmology.
@@
  • PBHs might be the supermassive or stupendously large BHs which seed the galaxy or structure formation Bean and Magueijo (2002); Kawasaki et al. (2012); Carr and Silk (2018); Carr et al. (2021). To satisfy the stringent constraints from CMB $\mu$- and $y$-distortion Kohri et al. (2014), large non-Gaussianities must be included Nakama et al. (2016, 2018, 2019); Atal et al. (2020b); Carr and Silk (2018); Liu and Bromm (2022); Biagetti et al. (2023); Gouttenoire et al. (2023); Hooper et al. (2023).
@@
  • According to the current observational constraints, asteroid-mass PBH can be all the dark matter, of which the accompanying induced GWs are detectable by space-borne interferometers LISA Barausse et al. (2020); Arun et al. (2022); Auclair et al. (2023); Bagui et al. (2023), Taiji Ren et al. (2023), and TianQin Liang et al. (2021). As positive non-Gaussianity can enhance the PBH abundance greatly, the required power spectrum of the curvature perturbation for $\Omega_{\mathrm{PBH}}=\Omega_{\mathrm{DM}}$ is reduced with positive non-Gaussianity, and the induced GW spectrum is also reduced, yet it is still above the LISA/Taiji/TianQin sensitivity curve Cai et al. (2019a); Bartolo et al. (2019); Unal (2019). The large non-Gaussian limit with $f_{\mathrm{NL}}\gg 1$ (i.e. the $\chi^{2}$-distribution) can be realized in the curvaton model when curvaton is negligible at its decay Bartolo et al. (2004); Enqvist and Nurmi (2005); Sasaki et al. (2006); Pi and Sasaki (2023a); Ferrante et al. (2022).
@@
  • In 2023, the Pulsar Timing Array collaborations NANOGrav Agazie et al. (2023a, b), EPTA combined with InPTA Antoniadis et al. (2023a, b, c), PPTA Zic et al. (2023); Reardon et al. (2023a, b), CPTA Xu et al. (2023), as well as IPTA Agazie et al. (2023c), announced the strong evidence of nHz stochastic GW background, while the induced GW interpretation is favored Afzal et al. (2023). However, as the amplitude of the nHz GW spectrum is relatively high ($\sim 10^{-9}$), sub-solar-mass PBHs might be overproduced, i.e. $\Omega_{\mathrm{PBH}}>\Omega_{\mathrm{DM}}$ This tension can be alleviated by introducing a negative non-Gaussianity with $f_{\mathrm{NL}}\lesssim-1$ to suppress the PBH formation Franciolini et al. (2023a); Liu et al. (2023); Choudhury et al. (2023a). Such a large and negative $f_{\mathrm{NL}}$ should be seen as the leading order of a fully nonlinear curvature perturbation, which can be realized in the curvaton scenario when the curvaton dominates at its decay moment Pi and Sasaki (2023a); Ferrante et al. (2022); Chen et al. (2023).
@@
  • It was shown that PBHs obey a Poisson distribution when the density perturbation is Gaussian Ali-Haïmoud (2018); Desjacques and Riotto (2018), while non-Gaussianity can make PBHs cluster initially on small scales Tada and Yokoyama (2015); Young and Byrnes (2015); Franciolini et al. (2018); Young and Byrnes (2020); Suyama and Yokoyama (2019); De Luca et al. (2021), which enhances the merger rate of PBHs significantly Clesse and García-Bellido (2017); Raidal et al. (2017). Such clustered PBHs can escape from the microlensing detections, which assume Poisson distribution García-Bellido and Clesse (2018).
@@
  • When the power spectrum of the curvature perturbation is enhanced on small scales, the loop corrections on CMB scales might be large, which may jeopardize the mechanism of enhancing the power spectrum on small scales to generate PBHs Kristiano and Yokoyama (2022, 2023). It was soon realized that the loop correction can be suppressed when the transition to slow-roll inflation is smooth Riotto (2023); Firouzjahi (2023); Firouzjahi and Riotto (2024); Firouzjahi (2024a). For recent discussions, see Choudhury et al. (2023b); Motohashi and Tada (2023); Tasinato (2023); Fumagalli (2023); Cheng et al. (2023); Motohashi and Tada (2023); Franciolini et al. (2023b); Fumagalli et al. (2023); Maity et al. (2023); Tada et al. (2024); Iacconi et al. (2023); Davies et al. (2024); Firouzjahi (2024b); Inomata (2024); Ballesteros and Egea (2024); Kristiano and Yokoyama (2024).
@@
For more discussion on the observational implications, see Part V (of the book Primordial Black Holes) and the references therein, for instance the review papers Carr et al. (2020); Carr and Kuhnel (2020); Carr et al. (2024).
@@
This paper is organized as follows. In Section II we review the calculation of PBH mass function, emphasizing on when and how the non-Gaussianity comes into play. In Section III, we study some typical inflation models which can generate large non-Gaussianity: the ultra-slow-roll inflation, constant-roll inflation with a bumpy potential, single-field inflation with piecewise quadratic potential, and the curvaton scenario. After getting the non-Gaussian curvature perturbations in these models, we calculate the PBH mass function and abundance by the Press-Schechter-type formalism in Section IV, as well as the scalar-induced GWs in Section V. We conclude by discussing some interesting topics to be studied in the future in Section VI.
@@ II.1 PBH abundance in the Gaussian case
In this section we will briefly review how to calculate the PBH mass function and abundance by the Press-Schechter formalism, emphasizing on where and how the non-linearities and primordial non-Gaussianities enter. Our starting point is the power spectrum of the primordial curvature perturbation on comoving slices, $\mathcal{P_{R}}(k)$, calculated at the end of inflation. In the simplest Press-Schechter formalism, the density contrast on comoving slices is used, which is related to $\mathcal{R}$ by the Poisson equation
$\delta\equiv\left(\frac{\delta\rho}{\rho_{0}}\right)_{c}\approx-\frac{2(1+w)}{5+3w}\frac{1}{H^{2}a^{2}}\nabla^{2}\mathcal{R}.$ (3)
@@
We will not discuss the issues of window function in this paper. The key point of the above calculation is that, if the curvature perturbation $\mathcal{R}$ is Gaussian, the linear approximation (3) tells us the density contrast is also Gaussian. Therefore the integration in (6) is simply a complementary error function, which gives a Gaussian suppression $\sim\exp(-\delta_{\mathrm{th}}^{2}/2\sigma_{\delta}^{2})$ in the far tail. This simple estimation is frequently quoted in the literature. However, it depends crucially on the Gaussianity of $\mathcal{R}$ and the linear approximation (3), both of which are doubtful when $\delta>\delta_{\mathrm{th}}$.